OpenStudy (anonymous):

6(2w+8)=4(5w-6) help solve?

OpenStudy (anonymous):

12w+48=25w+24 right?

OpenStudy (anonymous):

12w + 48 = 20w - 24 8w = 72 w = 9

OpenStudy (karatechopper):

distribute it first 12w+48=20w-24 then put the like terms together 12w-12w+48=20w-12w-24 48=8w-24 48+24=8w-24+24 72=8w 72/8=8w/8 9=w

OpenStudy (karatechopper):

@ayo u made 2 mistakes 1.u multiplied wrong twice

OpenStudy (anonymous):

yes i saw that haha